The base of an isosceles triangle is 8 feet kong. If the altitude is 6 feet long and increasing 3 inches per minute, at what rate are the base angles changing?


1
Expert's answer
2021-01-21T20:04:54-0500

Consider the isosceles triangle as shown in the figure below:





In the triangle,


"tan(\theta)=\frac{h}{4}"


Differentiate both sides with respect to "t" as,


"\frac{d}{dt}tan(\theta)=\frac{1}{4} \frac{d}{dt}(h)"


"sec^2(\theta)\frac{d\theta}{dt}=\frac{1}{4} \frac{dh}{dt}"


When "h=6,\frac{dh}{dt}=3" and "sec^2(\theta)=1+tan^2(\theta)=1+(\frac{6}{4})^2=\frac{13}{4}"


Therefore,


"\frac{13}{4}\frac{d\theta}{dt}=\frac{1}{4}(3)"


"13\frac{d\theta}{dt}=3"


"\frac{d\theta}{dt}=\frac{3}{13}"


Hence, the base angles are changing at rate "\frac{d\theta}{dt}=\frac{3}{13}" radian per minutes.